OpenStudy (anonymous):

I need help with an exam review question.. how do you write an equation for the line that is parallel to the given line and that passes through the given point? y=-5x+3; (-6,3)

OpenStudy (anonymous):

well when u have two line in parallel then m1=m2

OpenStudy (anonymous):

so in ur case the m1=m2=-5, and because the line is passing through the point (-6,3). then that yields to applying the following \[\frac{y-y_1}{x-x_1}=m\]

OpenStudy (anonymous):

sorry y1=5, x1=-6 and m=-5

OpenStudy (anonymous):

\[\frac{y-(6)}{x-(-3)}=-5\]

OpenStudy (anonymous):

do u know how to simplify this expression in the form y =mx+c

OpenStudy (anonymous):

no haha

OpenStudy (anonymous):

ok not to worry

OpenStudy (anonymous):

\[y-6=-5(x+3)\]

OpenStudy (anonymous):

oh wait .. so y=-5x+ ? idk

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

you're more than welcome all the best

OpenStudy (anonymous):

so I solve that?

OpenStudy (anonymous):

y-6=-5x-15

OpenStudy (anonymous):

y=-5x-9

OpenStudy (anonymous):

which is parallel to y=-5x+3

OpenStudy (anonymous):

thank you! this helps alot

OpenStudy (anonymous):

all the best...